need to understand how to get center and radius and then graph this equation?
x squared + y squared=1

Answer by jim_thompson5910(35256)

x^2 + y^2 = 1

can be rewritten as

(x-0)^2 + (y-0)^2 = 1^2

and that is in the form

(x-h)^2+(y-k)^2=r^2, which is a circle with radius r and center (h,k)

Since h = 0, k = 0 and r = 1, the center of x^2 + y^2 = 1 is (0,0) and it has a radius of 1 unit. This is a special circle known as the unit circle (which you'll use a lot in trigonometry).

The graph looks like this